# Long-Context Language Models?

### v2

@@ Abstract
The needle-in-a-haystack (NIAH) test, which examines the ability to retrieve a piece of information (the “needle”) from long distractor texts (the “haystack”), has been widely adopted to evaluate long-context language models (LMs). However, this simple retrieval-based test is indicative of only a superficial form of long-context understanding. To provide a more comprehensive evaluation of long-context LMs, we create a new synthetic benchmark Ruler with flexible configurations for customized sequence length and task complexity. Ruler expands upon the vanilla NIAH test to encompass variations with diverse types and quantities of needles. Moreover, Ruler introduces new task categories multi-hop tracing and aggregation to test behaviors beyond searching from context. We evaluate ten long-context LMs with 13 representative tasks in Ruler. Despite achieving nearly perfect accuracy in the vanilla NIAH test, all models exhibit large performance drops as the context length increases. While these models all claim context sizes of 32K tokens or greater, only four models (GPT-4, Command-R, Yi-34B, and Mixtral) can maintain satisfactory performance at the length of 32K. Our analysis of Yi-34B, which supports context length of 200K, reveals large room for improvement  as we increase input length and task complexity. We open source Ruler to spur comprehensive evaluation of long-context LMs. * Authors contributed equally.
@@ 1 Introduction
Using Ruler, we benchmark GPT-4  (OpenAI: Josh Achiam et al., 2023) and nine open-source models with context length ranging from 4k to 128k. Despite achieving nearly perfect performance on the vanilla NIAH test, all models exhibit large degradation on more complex tasks in Ruler as sequence length increases. While all models claim context size of 32k tokens or greater, our results indicate that only four of them can effectively handle sequence length of 32K by exceeding a qualitative threshold. Moreover, almost all models fall below the threshold before reaching the claimed context lengths. To obtain fine-grained model comparisons, we aggregate performance from 4k to 128k with two weighted average scores where the weights simulate the length distribution of real-world use cases. The top models - GPT-4, Command-R (Cohere, 2024), Yi-34B (Young et al., 2024), and Mixtral (Jiang et al., 2024), consistently outperform other models regardless of the chosen weighting scheme.
We further analyze Yi-34B, which claims context length of 200K and achieves the 2nd place on Ruler among open-source models. Our results demonstrate large degradation in Yi’s performance as we increase input length and task complexity. At large context sizes, Yi-34B often returns incomplete answers and fails to precisely locate the relevant information. Furthermore, we observe two behaviors emerging with the scaling of context size across multiple models: the increased reliance on parametric knowledge and the increased tendency to copy from context for non-retrieval tasks. Our additional ablations demonstrate that training on longer sequences does not always lead to better performance on Ruler, and that larger model sizes positively correlate with better long-context capabilities. Finally, we show that non-Transformer architectures, such as RWKV and Mamba, still lag behind Transformer by large margins on Ruler.
@@
  • We evaluate ten long-context LMs using Ruler and perform analysis across models and task complexities.
@@ 2 Related Work
Numerous long-context language models have been introduced lately owing to the progress in engineering, architectural, and algorithmic designs. Flash attention (Dao et al., 2022; Dao, 2023) and Ring attention (Liu et al., 2023) significantly reduce the memory footprint required for processing long context. Various sparse attention mechanisms (Child et al., 2019; Jaszczur et al., 2021) such as shifted sparse attention (Chen et al., 2024), dilated attention (Ding et al., 2023), and attention sinks (Han et al., 2023; Xiao et al., 2024b) were employed to enable efficient context scaling. Novel position embedding methods were proposed to improve length extrapolation in Transformers (Vaswani et al., 2017), including ALiBi (Press et al., 2022), xPOS (Sun et al., 2023b), and RoPE (Su et al., 2023) variants (Chen et al., 2023; Xiong et al., 2023; Peng et al., 2024; Liu et al., 2024b; Ding et al., 2024; Zhu et al., 2024). Another line of research focuses on reducing context size. This can be achieved by caching previous context using recurrence mechanism (Zhang et al., 2024a; Bulatov et al., 2023; Martins et al., 2022; Wu et al., 2022), or preserving only the salient information within long context via retrieval (Xu et al., 2024; Mohtashami & Jaggi, 2023; Wang et al., 2024; Tworkowski et al., 2024; Xiao et al., 2024a) or compression (Jiang et al., 2023). Finally, novel architectures (Gu et al., 2022; Fu et al., 2023a; Poli et al., 2023; Fu et al., 2023b; Sun et al., 2023a) such as Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) have also been proposed to efficiently handle long-context input.
@@
Our work is closely related to other works on benchmarking long-context language models. ZeroSCROLLS (Shaham et al., 2023) covers ten realistic natural language tasks, such as long-document QA and (query-based) summarization. L-Eval (An et al., 2024) also uses realistic data, which was filtered manually to ensure quality. LongBench (Bai et al., 2023) contains tasks in a bilingual setting. InfiniteBench (Zhang et al., 2024b) includes tasks with length greater than 100K tokens. LTM (Castillo et al., 2024) targets the evaluation of long-term conversations. To isolate the effect of parametric knowledge, previous works (Dong et al., 2023; Li et al., 2023b) also propose to use documents posted online later than a certain cutoff date, or leverage extremely low-resource materials (Tanzer et al., 2024). Compared to realistic benchmarks, synthetic tasks are more flexible to control the setup (e.g., sequence length and task complexity) and less affected by parametric knowledge. Recent works have mostly focused on retrieval-based synthetic tasks(Kamradt, 2023; Mohtashami & Jaggi, 2023; Li et al., 2023a; Liu et al., 2024c), with a few on other types of long-context usage, including various types of reasoning (Tay et al., 2021) and long-range discourse modeling (Sun et al., 2022).
@@ 3 The Ruler Benchmark
Ruler comprises tasks across four categories: retrieval, multi-hop tracing, aggregation, and question answering with all tasks configurable for varying length and complexity (see Table 2).
@@
(essays) ……
One of the special magic numbers for long-context is: 12345. ……
What is the special magic number for long-context mentioned in the provided text?
Answer: 12345
@@
(essays) ……
One of the special magic numbers for long-context is: 12345.
One of the special magic numbers for large-model is: 54321.
What is the special magic number for long-context mentioned in the provided text?
Answer: 12345
@@
(essays) ……
One of the special magic numbers for long-context is: 12345.
One of the special magic numbers for long-context is: 54321.
What are all the special magic numbers for long-context mentioned in the provided text?
Answer: 12345 54321
@@
(essays) ……
One of the special magic numbers for long-context is: 12345.
One of the special magic numbers for large-model is: 54321.
What are all the special magic numbers for long-context and large-model mentioned in the provided text?
Answer: 12345 54321
@@
(noises) ……
VAR X1 = 12345 …… VAR Y1 = 54321 ……
VAR X2 = X1 …… VAR Y2 = Y1 ……
VAR X3 = X2 …… VAR Y3 = Y2 ……
Find all variables that are assigned the value 12345.
Answer: X1 X2 X3
@@
aaa bbb ccc aaa ddd eee ccc fff ggg hhh iii iii ……
What are the 10 most common words in the above list?
Answer: aaa ccc iii ……
@@
aaa bbb ccc aaa ddd eee ccc fff ggg aaa hhh aaa ccc iii iii ……
What are the 3 most frequently appeared words in the above coded text?
Answer: aaa ccc iii
@@
Document 1: …… aaa ……
Document 2: …… bbb ……
Document 3: …… ccc ……
Question: question
Answer: bbb
@@ 3.4 Question Answering (QA)
Llama2-7B (chat) 4k - 85.6
GPT-4 128k 64k 96.6 96.3 95.2 93.2 87.0 81.2 91.6 89.0(1st) 94.1(1st)
Command-R (35B) 128k 32k 93.8 93.3 92.4 89.5 84.9 76.0 88.3 85.5(2nd) 91.1(2nd)
Yi (34B) 200k 32k 93.3 92.2 91.3 87.5 83.2 77.3 87.5 84.8(3th) 90.1(3th)
Mixtral (8x7B) 32k 32k 94.9 92.1 92.5 85.9 72.4 44.5 80.4 72.8(4th) 87.9(4th)
Mistral (7B) 32k 16k 93.6 91.2 87.2 75.4 49.0 13.8 68.4 55.6(7th) 81.2(5th)
ChatGLM (6B) 128k 4k 87.8 83.4 78.6 69.9 56.0 42.0 69.6 62.0(6th) 77.2(6th)
LWM (7B) 1M <4k 82.3 78.4 73.7 69.1 68.1 65.0 72.8 69.9(5th) 75.7(7th)
Together (7B) 32k 4k 88.2 81.1 69.4 63.0 0.0 0.0 50.3 33.8(8th) 66.7(8th)
LongChat (7B) 32k <4k 84.7 79.9 70.8 59.3 0.0 0.0 49.1 33.1(9th) 65.2(9th)
LongAlpaca (13B) 32k <4k 60.6 57.0 56.6 43.6 0.0 0.0 36.3 24.7(10th) 47.9(10th)
@@
4 Experiments & Results
Models & Inference setup
We select 10 long-context LLMs, including 9 open-source models and one closed-source model (GPT-4), covering diverse model sizes (6B to 8x7B with MoE architecture) and claimed context lengths (32k to 1M). Complete information about these models is included in Appendix A. We evaluate all models using vLLM (Kwon et al., 2023), an LLM serving system with efficient KV cache memory management. For all models, we run the inference in BFloat16 on 8 NVIDIA A100 GPUs with greedy decoding.
Task configurations
We test all models on 13 tasks ranging diverse complexities from the four categories of Ruler. The test configurations have been selected (shown in Appendix B) based on a task correlational study described in Appendix C. For each task, we evaluate each model with 500 examples generated for each length from the series (4k, 8k, 16k, 32k, 64k, 128k), while complying with each model’s necessary chat template. [^5] See Appendix D for model and tasks templates details. To prevent the model from refusing to answer a query or generating explanations, we append the task input with an answer prefix and check the presence of the target output with recall-based accuracy.
Effective Context Size
Model Ranking Criteria
@@ 4 Experiments & Results
We include the results of ten long-context LMs in comparison with the Llama2-7B baseline in Table 3.666Performance of base models and breakdown by task categories can be found in Appendix F. The performance at a certain length is the average of all 13 tasks in Ruler. While these models all claim effective context of 32K tokens or greater, none of them maintains performance above the Llama2-7B baseline at their claimed length, except for Mixtral, which achieves moderate performance on length doubling the claimed 32K context size. Despite achieving nearly perfect performance on the passkey retrieval and the vanilla NIAH task (shown in Appendix E), all models exhibit large degradation in RULER as sequence length increases. The best performant model on Ruler is GPT-4, which has the highest performance at length of 4k and demonstrates the least but non-marginal degradation (15.4) when extending the context to 128K. The top three ranked open-source models, Command-R, Yi-34B and Mixtral, all use a large base frequency in RoPE and are larger in parameter size than other models. Despite having been trained with context size of 1M, the LWM performs worse than Llama2-7B even at 4K. However, it shows smaller degradation with the increase of context size, therefore achieves higher rank than Mistral-7B when longer sequences receive larger weight (wAvg. inc). This result suggests a trade-off in evaluation between absolute performance on short sequences and the relative degradation with the scaling of context size.
@@ 5 Task Error Analysis
We evaluate Yi-34B-200K, the 2nd best open-source model on Ruler, with increased input lengths (up to 256K) on more complex tasks to understand the effect of task configurations and failure modes on Ruler.

### v3

@@ Abstract
The needle-in-a-haystack (NIAH) test, which examines the ability to retrieve a piece of information (the “needle”) from long distractor texts (the “haystack”), has been widely adopted to evaluate long-context language models (LMs). However, this simple retrieval-based test is indicative of only a superficial form of long-context understanding. To provide a more comprehensive evaluation of long-context LMs, we create a new synthetic benchmark Ruler with flexible configurations for customized sequence length and task complexity. Ruler expands upon the vanilla NIAH test to encompass variations with diverse types and quantities of needles. Moreover, Ruler introduces new task categories multi-hop tracing and aggregation to test behaviors beyond searching from context. We evaluate 17 long-context LMs with 13 representative tasks in Ruler. Despite achieving nearly perfect accuracy in the vanilla NIAH test, almost all models exhibit large performance drops as the context length increases. While these models all claim context sizes of 32K tokens or greater, only half of them can maintain satisfactory performance at the length of 32K. Our analysis of Yi-34B, which supports context length of 200K, reveals large room for improvement  as we increase input length and task complexity. We open source Ruler to spur comprehensive evaluation of long-context LMs. * Authors contributed equally.
@@ 1 Introduction
Using Ruler, we benchmark Gemini-1.5  (Reid et al., 2024), GPT-4  (OpenAI: Josh Achiam et al., 2023), and 15 open-source models with context length ranging from 4k to 128k. Despite achieving nearly perfect performance on the vanilla NIAH test, almost all models exhibit large degradation on more complex tasks in Ruler as sequence length increases. While all models claim context size of 32k tokens or greater, our results indicate that only half of them can effectively handle sequence length of 32K by exceeding a qualitative threshold. Moreover, almost all models fall below the threshold before reaching the claimed context lengths. To obtain fine-grained model comparisons, we aggregate performance from 4k to 128k with two weighted average scores where the weights simulate the length distribution of real-world use cases. The top two models - Gemini-1.5 and GPT-4, consistently outperform other models regardless of the chosen weighting scheme.
We further analyze Yi-34B, which claims context length of 200K and achieves reasonably good performance on Ruler among open-source models. Our results demonstrate large degradation in Yi’s performance as we increase input length and task complexity. At large context sizes, Yi-34B often returns incomplete answers and fails to precisely locate the relevant information. Furthermore, we observe two behaviors emerging with the scaling of context size across multiple models: the increased reliance on parametric knowledge and the increased tendency to copy from context for non-retrieval tasks. Our additional ablations demonstrate that training on longer sequences does not always lead to better performance on Ruler, and that larger model sizes positively correlate with better long-context capabilities. Finally, we show that non-Transformer architectures, such as RWKV and Mamba, still lag behind Transformer by large margins on Ruler.
@@
  • We evaluate 17 long-context LMs using Ruler and perform analysis across models and task complexities.
@@ 2 Related Work
Numerous long-context language models have been introduced lately owing to the progress in engineering, architectural, and algorithmic designs. Flash attention (Dao et al., 2022; Dao, 2023) and Ring attention (Liu et al., 2023) significantly reduce the memory footprint required for processing long context. Various sparse attention mechanisms (Child et al., 2019; Jaszczur et al., 2021) such as shifted sparse attention (Chen et al., 2024), dilated attention (Ding et al., 2023), and attention sinks (Han et al., 2023; Xiao et al., 2024b) were employed to enable efficient context scaling. Novel position embedding methods were proposed to improve length extrapolation in Transformers (Vaswani et al., 2017), including ALiBi (Press et al., 2022), xPOS (Sun et al., 2023b), and RoPE (Su et al., 2023) variants (Chen et al., 2023; Xiong et al., 2023; Peng et al., 2024; Liu et al., 2024b; Ding et al., 2024; Zhu et al., 2024). Another line of research focuses on reducing context size. This can be achieved by caching previous context using recurrence mechanism (Zhang et al., 2024a; Bulatov et al., 2023; Martins et al., 2022; Wu et al., 2022), retrieving relevant information from context (Xu et al., 2024a; Mohtashami & Jaggi, 2023; Wang et al., 2024; Tworkowski et al., 2024; Xiao et al., 2024a), or preserving the salient information via compression (Jiang et al., 2023). Finally, novel architectures (Gu et al., 2022; Fu et al., 2023a; Poli et al., 2023; Fu et al., 2023b; Sun et al., 2023a; Beck et al., 2024; Sun et al., 2024) such as Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) have also been proposed to efficiently handle long-context input.
@@
Our work is closely related to other works on benchmarking long-context language models. ZeroSCROLLS (Shaham et al., 2023) covers ten realistic natural language tasks, such as long-document QA and (query-based) summarization. L-Eval (An et al., 2024) also uses realistic data, which was filtered manually to ensure quality. LongBench (Bai et al., 2023) contains tasks in a bilingual setting. InfiniteBench (Zhang et al., 2024b) includes tasks with length greater than 100K tokens. LTM (Castillo et al., 2024) targets the evaluation of long-term conversations. To isolate the effect of parametric knowledge, previous works (Dong et al., 2023; Li et al., 2023b) also propose to use documents posted online later than a certain cutoff date, or leverage extremely low-resource materials (Tanzer et al., 2024). Compared to realistic benchmarks, synthetic tasks are more flexible to control the setup (e.g., sequence length and task complexity) and less affected by parametric knowledge. Recent works have primarily focused on retrieval-based synthetic tasks (Kamradt, 2023; Mohtashami & Jaggi, 2023; Li et al., 2023a; Liu et al., 2024d; Lee et al., 2024), with a few investigate other aspects, including fact reasoning (Kuratov et al., 2024; Karpinska et al., 2024), long-range discourse modeling (Sun et al., 2022), question answering (Levy et al., 2024; Yuan et al., 2024), many-shot in-context learning (Agarwal et al., 2024; Bertsch et al., 2024; Xu et al., 2024b), and code understanding (Liu et al., 2024c).
@@ 3 The Ruler Benchmark
Ruler comprises tasks across four categories: retrieval, multi-hop tracing, aggregation, and question answering. Evaluation examples in Ruler are automatically generated based on input configurations (see Table 2) that define the length and complexity of each input. Within a constrained domain as in RULER, the task complexity can be thought of as a function of the number of target output tokens and the signal-to-noise ratio in the context. We point readers to  (Goldman et al., 2024) for more comprehensive discussion on evaluation task design for long-context language models.
@@
(essays) …… One of the special magic numbers for long-context is: 12345. …… What is the special magic number for long-context mentioned in the provided text? Answer: 12345
@@
(essays) …… One of the special magic numbers for long-context is: 12345. One of the special magic numbers for large-model is: 54321. …… What is the special magic number for long-context mentioned in the provided text? Answer: 12345
@@
(essays) …… One of the special magic numbers for long-context is: 12345. One of the special magic numbers for long-context is: 54321. …… What are all the special magic numbers for long-context mentioned in the provided text? Answer: 12345 54321
@@
(essays) …… One of the special magic numbers for long-context is: 12345. One of the special magic numbers for large-model is: 54321. …… What are all the special magic numbers for long-context and large-model mentioned in the provided text? Answer: 12345 54321
@@
(noises) …… VAR X1 = 12345 …… VAR Y1 = 54321 …… VAR X2 = X1 …… VAR Y2 = Y1 …… VAR X3 = X2 …… VAR Y3 = Y2 …… Find all variables that are assigned the value 12345. Answer: X1 X2 X3
@@
aaa bbb ccc aaa ddd eee ccc fff ggg hhh iii iii …… What are the 10 most common words in the above list? Answer: aaa ccc iii ……
@@
aaa bbb ccc aaa ddd eee ccc fff ggg aaa hhh aaa ccc iii iii …… What are the 3 most frequently appeared words in the above coded text? Answer: aaa ccc iii
@@
Document 1: …… aaa …… Document 2: …… bbb …… Document 3: …… ccc …… Question: question Answer: bbb
@@ 3.4 Question Answering (QA)
4 Experiments & Results
Models & Inference setup
We select 17 long-context LLMs, including 15 open-source models and two closed-source model (Gemini-1.5-Pro and GPT-4 ), covering diverse model sizes (7B to 8x22B with MoE architecture) and claimed context lengths (32K to 1M). Complete information about these models is included in Appendix A. We evaluate all models using vLLM (Kwon et al., 2023), an LLM serving system with efficient KV cache memory management. For all models, we run the inference in BFloat16 on 8 NVIDIA A100 GPUs with greedy decoding.
Task configurations
We test all models on 13 tasks ranging diverse complexities from the four categories of Ruler. The test configurations have been selected (shown in Appendix B) based on a task correlational study described in Appendix C. We select these tasks as most models perform decently at short context size of 4K tokens. Our main goal is to see whether models can maintain such good performance with the scaling of context length. For each task, we evaluate each model with 500 examples generated for each length from the series (4K, 8K, 16K, 32K, 64K, 128K), while complying with each model’s necessary chat template. [^5] See Appendix D for model and tasks templates details. To prevent the model from refusing to answer a query or generating explanations, we append the task input with an answer prefix and check the presence of the target output with recall-based accuracy.
Effective Context Size
Model Ranking Criteria
@@ 4 Experiments & Results
Llama2 (7B) 4K - 85.6
Gemini-1.5-Pro 1M >128K 96.7 95.8 96.0 95.9 95.9 94.4 95.8 95.5(1st) 96.1(1st)
GPT-4 128K 64K 96.6 96.3 95.2 93.2 87.0 81.2 91.6 89.0(2nd) 94.1(2nd)
Llama3.1 (70B) 128K 64K 96.5 95.8 95.4 94.8 88.4 66.6 89.6 85.5(4th) 93.7(3rd)
Qwen2 (72B) 128K 32K 96.9 96.1 94.9 94.1 79.8 53.7 85.9 79.6(9th) 92.3(4th)
Command-R-plus (104B) 128K 32K 95.6 95.2 94.2 92.0 84.3 63.1 87.4 82.7(7th) 92.1(5th)
GLM4 (9B) 1M 64K 94.7 92.8 92.1 89.9 86.7 83.1 89.9 88.0(3rd) 91.7(6th)
Llama3.1 (8B) 128K 32K 95.5 93.8 91.6 87.4 84.7 77.0 88.3 85.4(5th) 91.3(7th)
GradientAI/Llama3 (70B) 1M 16K 95.1 94.4 90.8 85.4 80.9 72.1 86.5 82.6(8th) 90.3(8th)
Mixtral-8x22B (39B/141B) 64K 32K 95.6 94.9 93.4 90.9 84.7 31.7 81.9 73.5(11th) 90.3(9th)
Yi (34B) 200K 32K 93.3 92.2 91.3 87.5 83.2 77.3 87.5 84.8(6th) 90.1(10th)
Phi3-medium (14B) 128K 32K 93.3 93.2 91.1 86.8 78.6 46.1 81.5 74.8(10th) 88.3(11th)
Mistral-v0.2 (7B) 32K 16K 93.6 91.2 87.2 75.4 49.0 13.8 68.4 55.6(13th) 81.2(12th)
LWM (7B) 1M <4K 82.3 78.4 73.7 69.1 68.1 65.0 72.8 69.9(12th) 75.7(13th)
DBRX (36B/132B) 32K 8K 95.1 93.8 83.6 63.1 2.4 0.0 56.3 38.0(14th) 74.7(14th)
Together (7B) 32K 4K 88.2 81.1 69.4 63.0 0.0 0.0 50.3 33.8(15th) 66.7(15th)
LongChat (7B) 32K <4K 84.7 79.9 70.8 59.3 0.0 0.0 49.1 33.1(16th) 65.2(16th)
LongAlpaca (13B) 32K <4K 60.6 57.0 56.6 43.6 0.0 0.0 36.3 24.7(17th) 47.9(17th)
@@
We include the results of 17 long-context LMs in comparison with the Llama2-7B baseline in Table 3.666Performance of base models and breakdown by task categories can be found in Appendix F. The performance at a certain length is the average of all 13 tasks in Ruler. The closed-source model Gemini-1.5-Pro outperforms the rest of the models by a large margin, with the effective length greater than the maximum length we have tested on. Pressure testing this model with harder version of RULER can be interesting to follow up in the future. For the rest of the models, while they achieve nearly perfect performance on the passkey retrieval and the vanilla NIAH task (shown in Appendix E), all of them exhibit large degradation in RULER as sequence length increases and they fail to maintain performance above the Llama2-7B baseline at their claimed length. The top-ranked open-source models (Llama3.1, Qwen2 and Command-R-plus) share common configurations, such as having larger model sizes and using larger base frequencies in RoPE (Xiong et al., 2023). Large training context window is not always necessary for good long context performance – top-ranked open-source models contain both brute-force context scaling (Llama3.1 trained on 128K context length) and inference-time length extrapolation (Qwen2 trained on 32K context length). The less performant models also include those trained on much larger context size (e.g., LWM and GradientAI/Llama3 both on 1M context length). Although LWM achieves a higher rank than Mistral-v0.2 when longer sequences receive larger weight (wAvg. inc) and shows less degradation as the context size increases, it performs worse than Llama2-7B even at 4K. This result suggests a trade-off in evaluation between absolute performance on short sequences and the relative degradation with the scaling of context size. We provide more analysis on the model size and maximum training length in section 6.
@@ 5 Task Error Analysis
We evaluate Yi-34B-200K with increased input lengths (up to 256K) on more complex tasks to understand the effect of task configurations and failure modes on Ruler.